OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3) solve plz help

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

\[\huge{–2(x – 3) ≥ 5 – (x + 3)}\] \[\huge{-2x+6 \ge 5-x-3}\] \[\huge{-2x+x \ge 5-6-3}\] \[\huge{-x\ge -4}\] \[\huge{x \ge 4}\]

mathslover (mathslover):

the correct answer is \(\huge{x\le 4}\)

mathslover (mathslover):

sorry ..

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

thanks though